# Self-Normalized Resets for Plasticity in Continual Learning

### v1

@@ 4.3 Gradient Descent with Resets
$\frac{(1-p)^{T}}{(1-p)^{T}+1}\leq\frac{\lambda}{p}$
@@
$(1-p)^{T}\leq\frac{\frac{\lambda}{p}}{1-\frac{\lambda}{p}}=\frac{\lambda}{p-\lambda}$
@@
$(1-p)^{T}\leq\frac{\lambda}{p-\lambda}$
@@
Appendix B Learning a Single ReLU
In this section we prove Theorem 4.1 and Theorem 4.2. For convenience, we use slightly different notation from Section 4.1, such as treating the training examples $x$ as two-dimensional vectors with the element fixed as 1. For completeness, we begin by restating Section 4.1.
B.1 Preliminaries
@@
$f_{v}(x)=\sigma(v^{T}x)=\begin{cases}v^{T}x&\text{if $v^{T}x\geq 0$}\\ 0&\text{if $v^{T}x<0$}\end{cases}$
We refer to $v=(\tilde{v},b_{v})$ as the target parameters where $\tilde{v}$ is the slope and $b_{v}$ is the bias of the linear map $x\mapsto v^{T}x$. Likewise, we denote our model’s parameters as $w=(\tilde{w},b_{w})$ where $\tilde{w}$ is the slope and $b_{w}$ is the bias.
@@
$F(w)=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{T}x)-\sigma(v^{T}x))^{2}]$ (8)
@@
$\nabla F(w)=\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x]$ (9)
@@ B.1.1 L2 Init and L2 Regularization
$\bar{F}(w)=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{T}x)-\sigma(v^{T}x))^{2}]+\frac{\lambda}{2}||w-w_{0}||^{2}=F(w)+\frac{\lambda}{2}||w-w_{0}||^{2}$
@@
$\nabla F(w)=\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x]+\lambda(w-w_{0})=\nabla F(w)+\lambda(w-w_{0})$
@@
The proof follows immediately by Lemma B.4 for L2 Init. A trivial modification of Lemma B.4 yields an identical result for L2 regularization, which we omit for brevity. ∎
B.1.2 Gradient Descent with Resets
@@
$\mathcal{O}_{\epsilon}(w)=\begin{cases}\text{True}&\text{ if }\sup_{x\in[-L,L]\times\{1\}}w^{T}x\leq\epsilon\\ \text{False}&\text{ if }\sup_{x\in[-L,L]\times\{1\}}w^{T}x>\epsilon\end{cases}$
@@
The theorem is restated and proven as Theorem B.3. ∎
B.2 Proofs
B.2.1 Properties of the Loss Function
Lemma B.1.
@@ B.2.1 Properties of the Loss Function
Lemma B.2.
Suppose that $x$ is sampled according to $x\sim$ Uniform$(-L,L)\times\{1\}\subseteq\mathbb{R}^{2}$ and that the target parameters $v=(\tilde{v},b_{v})$ satisfy $\tilde{v}>0$ and $b_{v}\leq 0$. Then for $w\leq 0$, $\nabla F(w)=\mathbb{E}_{x}[\sigma(w^{T}x)x]$.
@@
For almost every $\tilde{x}>0$, $w^{T}x=\tilde{w}\tilde{x}+b_{w}<0$ and so
$(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x$ $=0$ by $\mathbb{I}(w^{T}x\geq 0)=0$
$=\sigma(w^{T}x)x$ by $\sigma(w^{T}x)=0$ since $w^{T}x<0$
As for $\tilde{x}<0$, we have that $v^{T}x<0$ since $\tilde{v}>0$ and $b_{v}\leq 0$, and so $\sigma(v^{T}x)=0$. Hence,
$(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x=\sigma(w^{T}x)\mathbb{I}(w^{T}x\geq 0)x=\sigma(w^{T}x)x$
where the last equality follows by the fact that $\mathbb{I}(w^{T}x\geq 0)$ is redundant given that $\sigma(w^{T}x)=0$ if $w^{T}x<0$. Therefore, for almost every $x\in\text{ Uniform}(-L,L)\times\{1\}$, we have that
$(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x=\sigma(w^{T}x)x$ (15)
@@
$\nabla F(w)=\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x]$
@@
$\nabla F(w)=\mathbb{E}_{x}[\sigma(w^{T}x)x]$
Lemma B.3.
@@
By $w,u\leq 0$ and Lemma B.2 we have that $\nabla F(w)=\mathbb{E}_{x}[\sigma(w^{T}x)x]$ and $\nabla F(u)=\mathbb{E}_{x}[\sigma(u^{T}x)x]$. Then we establish the desired result as follows,
$||\nabla F(w)-\nabla F(u)||$ $=||\mathbb{E}_{x}[\sigma(w^{T}x)x-\sigma(u^{T}x)x]||$
$\leq\mathbb{E}_{x}[||\sigma(w^{T}x)x-\sigma(u^{T}x)x||]$ by Jensen’s inequality
$=\mathbb{E}_{x}[|\sigma(w^{T}x)-\sigma(u^{T}x)|\cdot||x||]$
$\leq\mathbb{E}_{x}[|w^{T}x-u^{T}x|\cdot||x||]$ by Lemma B.1
@@
B.2.2 Convergence After a Negative Initialization
Lemma B.4.
@@ B.2.2 Convergence After a Negative Initialization
$=(1-\eta\lambda)b_{w_{t}}-\eta\mathbb{E}_{x}[\sigma(w^{T}_{t}x)]$ by Lemma B.2 given that $w_{t}\leq 0,\tilde{v}>0,b_{v}\leq 0$
$\leq(1-\eta\lambda)b_{w_{t}}$ since $\sigma(w_{t}^{T}x)\geq 0,\forall x$
@@
Next, for (16) we have that $\tilde{w}_{t+1}\leq 0$ if $-L\tilde{w}_{t+1}+b_{w_{t+1}}=w_{t+1}^{T}(-L,1)\geq 0$ since $b_{w_{t+1}}\leq 0$. This is equivalent to showing that
$(w_{t}-\eta\nabla F(w_{t})-\eta\lambda(w_{t}-w_{0}))^{T}(-L,1)\geq 0$ (20)
We proceed by bounding $\eta\nabla F(w_{t})^{T}(-L,1)$. Again invoking Lemma B.2, we have that
$\eta\nabla F(w_{t})^{T}(-L,1)$ $=\eta\mathbb{E}_{x}[\sigma(w_{t}^{T}x)x]^{T}(-L,1)$
$=\eta\mathbb{E}_{x}[\sigma(w_{t}^{T}x)(-L\tilde{x}+1)]$
$\leq\eta\sigma(w_{t}^{T}(-L,1))(L^{2}+1)$ by assumption that $\tilde{w}_{t}\leq 0$
$=\alpha\sigma(w_{t}^{T}(-L,1))$ by choice of $\eta$
$=\alpha w^{T}_{t}(-L,1)$ by assumption of the invariant (19)
$\leq w^{T}_{t}(-L,1)$ since $\alpha\leq 1$
@@
$(w_{t}-\eta\nabla F(w_{t})-\eta\lambda(w_{t}-w_{0}))^{T}(-L,1)$ $\geq(1-\eta\lambda)w_{t}^{T}(-L,1)+\eta\lambda w_{0}^{T}(-L,1)-w_{t}^{T}(-L,1)$
$\geq(1-\eta\lambda)w_{t}^{T}(-L,1)+\eta\lambda w_{t}^{T}(-L,1)-w_{t}^{T}(-L,1)$
@@
$=(1-\eta\lambda)\tilde{w_{t}}+\eta\lambda\tilde{w}_{0}-\eta\mathbb{E}_{x}[\sigma(w^{T}_{t}x)\tilde{x}]$ by Lemma B.2 given that $w_{t}\leq 0,\tilde{v}>0,b_{v}\leq 0$
$\geq(1-\eta\lambda)\tilde{w_{t}}+\eta\lambda\tilde{w}_{0}$ since $\tilde{x}>0\Rightarrow\sigma(w_{t}^{T}x)=0$ by $w_{t}\leq 0$
@@
$F(w_{t})$ $=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{T}_{t}x)-\sigma(v^{T}x))^{2}]$
$\geq\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{T}_{t}x)-\sigma(v^{T}x))^{2}\mathbb{I}_{\{\tilde{x}\geq 0\}}]$
$=\mathbb{E}_{x}[\frac{1}{2}(\sigma(v^{T}x))^{2}\mathbb{I}_{\{\tilde{x}\geq 0\}}]$ since $w_{t}\leq 0$ and hence $\sigma(w_{t}^{T}x)\mathbb{I}_{\{\tilde{x}\geq 0\}}=0$
$=\mathbb{E}_{x}[\frac{1}{2}(\sigma(v^{T}x))^{2}]$ since $v_{t}^{T}x\leq 0$ for $\tilde{x}<0$ by $\tilde{v}>0$, $b_{v}\leq 0$
@@
B.2.3 Resets After a Negative Initialization
Lemma B.5.
Suppose that $x$ is sampled according to $x\sim$ Uniform$(-L,L)\times\{1\}\subseteq\mathbb{R}^{2}$ and that the target parameters $v=(\tilde{v},b_{v})$ satisfy $\tilde{v}>0$ and $b_{v}\leq 0$. Then applying gradient descent according to equation (10) with $w_{0}$ satisfying $\tilde{w_{0}}\in[-1,0]$ and $b_{w_{0}}=0$ and learning rate $\eta<\frac{1}{L^{2}+1}$, we have that for any threshold $\epsilon>0$, there exists some $t\in\{0,1,\ldots,T=\lceil\frac{16L^{2}C_{\eta}}{\epsilon^{4}}\rceil\}$, where $C_{\eta}=\frac{2}{\eta}F((-1,0))$, such that $\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{T}x)\leq\epsilon$.
@@ B.2.3 Resets After a Negative Initialization
$\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{T}x)=-L\tilde{w}_{0}\leq\epsilon$
and so at $t=0$ we immediately have that $\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{T}x)\leq\epsilon$. Therefore, we consider the case of $\tilde{w}_{0}\in[-L,-\frac{\epsilon}{L})$ and proceed as follows. According to Lemma B.4 and by the choice of learning rate $\eta$, we have that $\forall t\in\{0,1,\ldots,T\},w_{t}\leq 0$. Therefore by Lemma B.2, $\nabla F(\cdot)$ is $(L^{2}+1)$-Lipschitz continuous on the span of $\{w_{0},w_{1},\ldots,w_{T}\}$. Consequently, following the canonical analysis of gradient descent, we have that for any $t\geq 0$
$F(w_{t+1})$ $\leq F(w_{t})+\nabla F(w_{t})^{T}(w_{t+1}-w_{t})+\frac{L^{2}+1}{2}||w_{t+1}-w_{t}||^{2}$
@@
$||\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)x]||^{2}\leq\frac{C_{\eta}}{T}$
@@
$\mathbb{E}_{x}[(\sigma(w_{t}^{T}x)-\sigma(v^{T}x))\mathbb{I}(w^{T}x\geq 0)]^{2}\leq\frac{C_{\eta}}{T}$
Additionally, by Lemma B.4 we have that $\tilde{w}_{0}\leq\tilde{w}_{t}\leq 0$ and $b_{w_{t}}\leq 0$, and thus by Lemma B.2, we have that
$\mathbb{E}_{x}[\sigma(w^{T}x)]^{2}=\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))\mathbb{I}(w_{t}^{T}x\geq 0)]^{2}\leq\frac{C_{\eta}}{T}$
@@
$\mathbb{E}_{x}[\sigma(w_{t}^{T}x)]\leq\sqrt{\frac{C_{\eta}}{T}}$
Then, noting that $\tilde{w}_{0}\leq\tilde{w}_{t}\leq 0$ and $b_{w_{t}}\leq 0$, we have that $2L\mathbb{E}_{x}[\sigma(w_{t}^{T}x)]$ is the area of the triangle formed by the line $w_{t}^{T}x$ over the $x$-axis with its base ranging from its $x$-intercept to $-L$. We denote the length of its base by $b$ and its height by $h$. Given that $\tilde{w}_{t}\leq 0$ it follows that $h=\sigma(w_{t}^{T}(-L,1))=\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{T}x)$. Additionally, we have that $h=|\tilde{w}_{t}|b\leq|\tilde{w}_{0}|b\leq b$ given that $-1\leq\tilde{w}_{0}\leq\tilde{w}_{t}\leq 0$. Hence we have that
$\frac{h^{2}}{4L}\leq\frac{bh}{4L}=\mathbb{E}_{x}[\sigma(w_{t}^{T}x)]\leq\sqrt{\frac{C_{\eta}}{T}}$
@@
$\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{T}x)=h\leq\sqrt{4L\sqrt{\frac{C_{\eta}}{T}}}\leq\epsilon$
B.2.4 Convergence After a Positive Initialization
Lemma B.6.

### v2

@@ Appendix C Additional Experimental Details and Hyperparameter Sweep
$\frac{(1-p)^{\top}}{(1-p)^{\top}+1}\leq\frac{\lambda}{p}$
@@
$(1-p)^{\top}\leq\frac{\frac{\lambda}{p}}{1-\frac{\lambda}{p}}=\frac{\lambda}{p-\lambda}$
@@
$(1-p)^{\top}\leq\frac{\lambda}{p-\lambda}$
@@
Proposition D.2.
Restatement of Proposition 2.2] The ratio of total error rate with a fixed threshold $r^{*}$ to that under SNR scales like
$\Omega\left(\exp\left(\log(\alpha(1-\alpha)^{-1})\left(-\frac{1}{2}+\frac{1}{2}\frac{\log(1-p_{1})}{\log(1-p_{2})}\right)\right)\right)$
Proof.
For notational convenience, define $\bar{\alpha}=\alpha(1-\alpha)^{-1}$. Notice that by Proposition 2.1, under SNR, neuron $i$, $(I=1,2)$, is reset if it inactive for any longer that time $\bar{T}=\log(\bar{\alpha}))/\log(1-p_{i})$. Consequently, the expected delay penalty, $\mathbb{E}[\tau|H_{0}]+\mathbb{E}[\tau|H_{1}]$ for neuron $i$, is simply
$\frac{\log(\bar{\alpha})}{\log(1-p_{i})}+\frac{1}{p_{i}}(1-\bar{\alpha})$
Letting the optimal fixed threshold be $r^{*}$, we must have that the expected total delay across both neurons under this fixed threshold is at least $2r^{*}$. This total expected delay can be no larger than that under SNR. Thus,
$2r^{*}\leq\log(\bar{\alpha})\left(\frac{1}{\log(1-p_{1})}+\frac{1}{\log(1-p_{2})}\right)+(1-\bar{\alpha})\left(\frac{1}{p_{1}}+\frac{1}{p_{2}}\right)$
But the sum of the error rates across the two neurons with the fixed threshold $r^{*}$ is at least $(1-p_{1})^{r^{*}}+\bar{\alpha}$, while the total error rate under SNR is precisely $\bar{\alpha}$. Dividing these two quantities and employing the upper bound derived on $r^{*}$ then yields the result. ∎
Appendix E Learning a Single ReLU
In this section we prove Theorem 4.1.
E.1 Preliminaries
@@
$f_{v}(x)=\sigma(v^{\top}x)=\begin{cases}v^{\top}x&\text{if $v^{\top}x\geq 0$}\\ 0&\text{if $v^{\top}x<0$}\end{cases}$
We refer to $v=(\tilde{v},b_{v})$ as the target parameters where $\tilde{v}$ is the slope and $b_{v}$ is the bias of the linear map $x\mapsto v^{\top}x$. Likewise, we denote our model’s parameters as $w=(\tilde{w},b_{w})$ where $\tilde{w}$ is the slope and $b_{w}$ is the bias.
@@
$F(w)=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{\top}x)-\sigma(v^{\top}x))^{2}]$ (3)
@@
$\nabla F(w)=\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x]$ (4)
@@ E.1.1 L2 Init and L2 Regularization
$\bar{F}(w)=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{\top}x)-\sigma(v^{\top}x))^{2}]+\frac{\lambda}{2}||w-w_{0}||^{2}=F(w)+\frac{\lambda}{2}||w-w_{0}||^{2}$
@@
$\nabla F(w)=\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x]+\lambda(w-w_{0})=\nabla F(w)+\lambda(w-w_{0})$
@@
The proof follows immediately by Lemma E.4 for L2 Init. A trivial modification of Lemma E.4 yields an identical result for L2 regularization, which we omit for brevity. ∎
E.1.2 Gradient Descent with Resets
@@
$\mathcal{O}_{\epsilon}(w)=\begin{cases}\text{True}&\text{ if }\sup_{x\in[-L,L]\times\{1\}}w^{\top}x\leq\epsilon\\ \text{False}&\text{ if }\sup_{x\in[-L,L]\times\{1\}}w^{\top}x>\epsilon\end{cases}$
@@
The theorem is restated and proven as Theorem E.3. ∎
E.2 Proofs
E.2.1 Properties of the Loss Function
Lemma E.1.
@@ E.2.1 Properties of the Loss Function
Lemma E.2.
Suppose that $x$ is sampled according to $x\sim$ Uniform$(-L,L)\times\{1\}\subseteq\mathbb{R}^{2}$ and that the target parameters $v=(\tilde{v},b_{v})$ satisfy $\tilde{v}>0$ and $b_{v}\leq 0$. Then for $w\leq 0$, $\nabla F(w)=\mathbb{E}_{x}[\sigma(w^{\top}x)x]$.
@@
For almost every $\tilde{x}>0$, $w^{\top}x=\tilde{w}\tilde{x}+b_{w}<0$ and so
$(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x$ $=0$ by $\mathbb{I}(w^{\top}x\geq 0)=0$
$=\sigma(w^{\top}x)x$ by $\sigma(w^{\top}x)=0$ since $w^{\top}x<0$
As for $\tilde{x}<0$, we have that $v^{\top}x<0$ since $\tilde{v}>0$ and $b_{v}\leq 0$, and so $\sigma(v^{\top}x)=0$. Hence,
$(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x=\sigma(w^{\top}x)\mathbb{I}(w^{\top}x\geq 0)x=\sigma(w^{\top}x)x$
where the last equality follows by the fact that $\mathbb{I}(w^{\top}x\geq 0)$ is redundant given that $\sigma(w^{\top}x)=0$ if $w^{\top}x<0$. Therefore, for almost every $x\in\text{ Uniform}(-L,L)\times\{1\}$, we have that
$(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x=\sigma(w^{\top}x)x$ (10)
@@
$\nabla F(w)=\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x]$
@@
$\nabla F(w)=\mathbb{E}_{x}[\sigma(w^{\top}x)x]$
Lemma E.3.
@@
By $w,u\leq 0$ and Lemma E.2 we have that $\nabla F(w)=\mathbb{E}_{x}[\sigma(w^{\top}x)x]$ and $\nabla F(u)=\mathbb{E}_{x}[\sigma(u^{\top}x)x]$. Then we establish the desired result as follows,
$||\nabla F(w)-\nabla F(u)||$ $=||\mathbb{E}_{x}[\sigma(w^{\top}x)x-\sigma(u^{\top}x)x]||$
$\leq\mathbb{E}_{x}[||\sigma(w^{\top}x)x-\sigma(u^{\top}x)x||]$ by Jensen’s inequality
$=\mathbb{E}_{x}[|\sigma(w^{\top}x)-\sigma(u^{\top}x)|\cdot||x||]$
$\leq\mathbb{E}_{x}[|w^{\top}x-u^{\top}x|\cdot||x||]$ by Lemma E.1
@@
E.2.2 Convergence After a Negative Initialization
Lemma E.4.
@@ E.2.2 Convergence After a Negative Initialization
$=(1-\eta\lambda)b_{w_{t}}-\eta\mathbb{E}_{x}[\sigma(w^{\top}_{t}x)]$ by Lemma E.2 given that $w_{t}\leq 0,\tilde{v}>0,b_{v}\leq 0$
$\leq(1-\eta\lambda)b_{w_{t}}$ since $\sigma(w_{t}^{\top}x)\geq 0,\forall x$
@@
Next, for (11) we have that $\tilde{w}_{t+1}\leq 0$ if $-L\tilde{w}_{t+1}+b_{w_{t+1}}=w_{t+1}^{\top}(-L,1)\geq 0$ since $b_{w_{t+1}}\leq 0$. This is equivalent to showing that
$(w_{t}-\eta\nabla F(w_{t})-\eta\lambda(w_{t}-w_{0}))^{\top}(-L,1)\geq 0$ (15)
We proceed by bounding $\eta\nabla F(w_{t})^{\top}(-L,1)$. Again invoking Lemma E.2, we have that
$\eta\nabla F(w_{t})^{\top}(-L,1)$ $=\eta\mathbb{E}_{x}[\sigma(w_{t}^{\top}x)x]^{\top}(-L,1)$
$=\eta\mathbb{E}_{x}[\sigma(w_{t}^{\top}x)(-L\tilde{x}+1)]$
$\leq\eta\sigma(w_{t}^{\top}(-L,1))(L^{2}+1)$ by assumption that $\tilde{w}_{t}\leq 0$
$=\alpha\sigma(w_{t}^{\top}(-L,1))$ by choice of $\eta$
$=\alpha w^{\top}_{t}(-L,1)$ by assumption of the invariant (14)
$\leq w^{\top}_{t}(-L,1)$ since $\alpha\leq 1$
@@
$(w_{t}-\eta\nabla F(w_{t})-\eta\lambda(w_{t}-w_{0}))^{\top}(-L,1)$ $\geq(1-\eta\lambda)w_{t}^{\top}(-L,1)+\eta\lambda w_{0}^{\top}(-L,1)-w_{t}^{\top}(-L,1)$
$\geq(1-\eta\lambda)w_{t}^{\top}(-L,1)+\eta\lambda w_{t}^{\top}(-L,1)-w_{t}^{\top}(-L,1)$
@@
$=(1-\eta\lambda)\tilde{w_{t}}+\eta\lambda\tilde{w}_{0}-\eta\mathbb{E}_{x}[\sigma(w^{\top}_{t}x)\tilde{x}]$ by Lemma E.2 given that $w_{t}\leq 0,\tilde{v}>0,b_{v}\leq 0$
$\geq(1-\eta\lambda)\tilde{w_{t}}+\eta\lambda\tilde{w}_{0}$ since $\tilde{x}>0\Rightarrow\sigma(w_{t}^{\top}x)=0$ by $w_{t}\leq 0$
@@
$F(w_{t})$ $=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{\top}_{t}x)-\sigma(v^{\top}x))^{2}]$
$\geq\mathbb{E}_{x}[\frac{1}{2}(\sigma(w^{\top}_{t}x)-\sigma(v^{\top}x))^{2}\mathbb{I}_{\{\tilde{x}\geq 0\}}]$
$=\mathbb{E}_{x}[\frac{1}{2}(\sigma(v^{\top}x))^{2}\mathbb{I}_{\{\tilde{x}\geq 0\}}]$ since $w_{t}\leq 0$ and hence $\sigma(w_{t}^{\top}x)\mathbb{I}_{\{\tilde{x}\geq 0\}}=0$
$=\mathbb{E}_{x}[\frac{1}{2}(\sigma(v^{\top}x))^{2}]$ since $v_{t}^{\top}x\leq 0$ for $\tilde{x}<0$ by $\tilde{v}>0$, $b_{v}\leq 0$
@@
E.2.3 Resets After a Negative Initialization
Lemma E.5.
Suppose that $x$ is sampled according to $x\sim$ Uniform$(-L,L)\times\{1\}\subseteq\mathbb{R}^{2}$ and that the target parameters $v=(\tilde{v},b_{v})$ satisfy $\tilde{v}>0$ and $b_{v}\leq 0$. Then applying gradient descent according to equation (5) with $w_{0}$ satisfying $\tilde{w_{0}}\in[-1,0]$ and $b_{w_{0}}=0$ and learning rate $\eta<\frac{1}{L^{2}+1}$, we have that for any threshold $\epsilon>0$, there exists some $t\in\{0,1,\ldots,T=\lceil\frac{16L^{2}C_{\eta}}{\epsilon^{4}}\rceil\}$, where $C_{\eta}=\frac{2}{\eta}F((-1,0))$, such that $\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{\top}x)\leq\epsilon$.
@@ E.2.3 Resets After a Negative Initialization
$\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{\top}x)=-L\tilde{w}_{0}\leq\epsilon$
and so at $t=0$ we immediately have that $\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{\top}x)\leq\epsilon$. Therefore, we consider the case of $\tilde{w}_{0}\in[-L,-\frac{\epsilon}{L})$ and proceed as follows. According to Lemma E.4 and by the choice of learning rate $\eta$, we have that $\forall t\in\{0,1,\ldots,T\},w_{t}\leq 0$. Therefore by Lemma E.2, $\nabla F(\cdot)$ is $(L^{2}+1)$-Lipschitz continuous on the span of $\{w_{0},w_{1},\ldots,w_{T}\}$. Consequently, following the canonical analysis of gradient descent, we have that for any $t\geq 0$
$F(w_{t+1})$ $\leq F(w_{t})+\nabla F(w_{t})^{\top}(w_{t+1}-w_{t})+\frac{L^{2}+1}{2}||w_{t+1}-w_{t}||^{2}$
@@
$||\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)x]||^{2}\leq\frac{C_{\eta}}{T}$
@@
$\mathbb{E}_{x}[(\sigma(w_{t}^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w^{\top}x\geq 0)]^{2}\leq\frac{C_{\eta}}{T}$
Additionally, by Lemma E.4 we have that $\tilde{w}_{0}\leq\tilde{w}_{t}\leq 0$ and $b_{w_{t}}\leq 0$, and thus by Lemma E.2, we have that
$\mathbb{E}_{x}[\sigma(w^{\top}x)]^{2}=\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))\mathbb{I}(w_{t}^{\top}x\geq 0)]^{2}\leq\frac{C_{\eta}}{T}$
@@
$\mathbb{E}_{x}[\sigma(w_{t}^{\top}x)]\leq\sqrt{\frac{C_{\eta}}{T}}$
Then, noting that $\tilde{w}_{0}\leq\tilde{w}_{t}\leq 0$ and $b_{w_{t}}\leq 0$, we have that $2L\mathbb{E}_{x}[\sigma(w_{t}^{\top}x)]$ is the area of the triangle formed by the line $w_{t}^{\top}x$ over the $x$-axis with its base ranging from its $x$-intercept to $-L$. We denote the length of its base by $b$ and its height by $h$. Given that $\tilde{w}_{t}\leq 0$ it follows that $h=\sigma(w_{t}^{\top}(-L,1))=\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{\top}x)$. Additionally, we have that $h=|\tilde{w}_{t}|b\leq|\tilde{w}_{0}|b\leq b$ given that $-1\leq\tilde{w}_{0}\leq\tilde{w}_{t}\leq 0$. Hence we have that
$\frac{h^{2}}{4L}\leq\frac{bh}{4L}=\mathbb{E}_{x}[\sigma(w_{t}^{\top}x)]\leq\sqrt{\frac{C_{\eta}}{T}}$
@@
$\sup_{x\in[-L,L]\times\{1\}}\sigma(w_{t}^{\top}x)=h\leq\sqrt{4L\sqrt{\frac{C_{\eta}}{T}}}\leq\epsilon$
E.2.4 Convergence After a Positive Initialization
Lemma E.6.